Find the sum, (if it exists), of the infinite geometric sequence with first term 2 and r = 1/3
thanks for the help! i've been stuck on this one for so long!

Answer by sudhanshu_kmr(1152)


it means.....
2, 2/3 , 2/9, ............infinity
sum = a/(1-r)
= 2/(1-1/3)
= 2/(2/3)
= 3